7) The path of a falling object is given by the function where represents the initial velocity in ft/sec and represents the initial height.
a) If a rock is thrown upward with an initial velocity of 32 feet per second from the top of a 40-foot building, write the height equation using this information.

b) How high is the rock after 0.5 seconds?


c) After how many seconds will the rock reach maximum height?

d) What is the maximum height?

Answer by stanbon(75887)

7) The path of a falling object is given by the function where represents the initial velocity in ft/sec and represents the initial height.
a) If a rock is thrown upward with an initial velocity of 32 feet per second from the top of a 40-foot building, write the height equation using this information.
h(t) = -16t^2+32t+40
b) How high is the rock after 0.5 seconds?
h(0.5) = -16*(0.5)^2+32*0.5+40
h(0.5) = 52 ft.
---------------
c) After how many seconds will the rock reach maximum height?
Maximum occurs at t=-b/2a = -32/(2*-16) = 1 second
---------------
d) What is the maximum height?
h(1) = -16+32+40 = 56
===============
Cheers,
Stan H.